what is the vertex form to 3x^2+12x+20
Found 2 solutions by jim_thompson5910, Onmyoji.S:
Answer by jim_thompson5910(35256)

y = 3x^2+12x+20

y - 20 = 3x^2+12x

y - 20 = 3(x^2+4x)

y - 20 = 3(x^2+4x+4-4)

y - 20 = 3((x^2+4x+4)-4)

y - 20 = 3((x+2)^2-4)

y - 20 = 3(x+2)^2-3*4

y - 20 = 3(x+2)^2-12

y = 3(x+2)^2-12+20

y = 3(x+2)^2 + 8

===================================================================

So after completing the square,

3x^2+12x+20

turns into

3(x+2)^2 + 8

which is now in vertex form.

Answer by Onmyoji.S(32)

3x²+12x+20
because x-coordinate of the vertex form is -b/2a
so x=-12/2(3)
x=2